OpenStudy (annie96):

Select the pair of equations whose graphs are perpendicular. (Points : 5) 5x – 8y = 9 12x – 5y = 7 3x + 6y = 8 y = 3x – 8 y = 2x – 7 x + 3y = 3 2y = –3x + 5 2x – 3y = 4 Combine like terms to simplify the expression: 7x + 3y – 2 + 6x + 1 + y2 (Points : 5) y2 + 3y + 13x – 1 3y3 + 13x2 + 1 2y3 + 13x – 1 y2 + 3y + 13x2 + 1 Given the linear equation 3x + y = 5, find the slope of its graph. (Points : 5) –3 –5 3 5

OpenStudy (annie96):

Solve the system. x – 2y = 10 3x + 2y = 14 (Points : 5) (–8, 6) (–2, 6) (6, –8) (6, –2)

OpenStudy (anonymous):

the sum of two consecutive odd numbers means a+b where, because they are consecutive odd number they have a difference of two, so b = a+2 so the sum of two consecutive numbers will be a + a + 2 = S (a was the first number so it is the smallest) S=a+a+2 S=2a+2 you can do it from there 6.6 I can't see the picture...I don't know what the question is asking (although the answers suggest they are looking for the union which is what set of numbers do both sets share. 8.8 lines are perpendicular or when they are oriented 90 degrees apart the product of their respective slopes will be negative one so just find the pair where m1*m2=-1 10.10 y=dy/dx*x+(y1-dy/dx*x1) 11.11 x+y=5 rearrange in standard point-slope form (y=mx+b) and the slope will be m or the leading coefficient in front of x 12.12 rearrange in standard form and solve for y when x equals zero, this is the y intercept 16.16 the y's cancel out just by adding the two equations together, solve for x and then plug that value of x into one of the original equations to solve for y. i dont know if this is right or not

OpenStudy (anonymous):

is there an answer choice

OpenStudy (anonymous):

Change to y = mx + b format 2y = -3x + 5 Divide EVERYTHING by 2y to isolate y y = -3/2x + 5/2 2x - 3y = 4 Subtract 2x from both sides -3y = -2x + 4 Divide EVERYTHING by -3 y = 2/3x + 4 2y = -3x + 5 2x -3y = 4 These two equations are perpendicular because -3/2 is the reciprocal of 2/3

OpenStudy (anonymous):

i guess i dont know

OpenStudy (anonymous):

please reply if i got this correct